# Fractional quantum Hall effect of partons and the nature of the 8/17 state in the zeroth Landau level of bilayer graphene

### v1

@@ Abstract
The Haldane-Halperin hierarchy postulates that the quasiholes or quasiparticle excitations of a parent FQHE state condense into a daughter FQHE state and this scheme can produce candidate states for any odd-denominator fraction [4, 5]. Jain’s CF theory builds many-particle correlations by the process of vortex attachment wherein electrons bind zeros of the many-body wave function to turn into CFs which to the zeroth-order approximation are taken to be non-interacting [6]. A vast majority of the FQHE phenomenology, especially that observed in the lowest Landau level (LLL), is well-described by the CF theory. In particular, the CF theory explains why the strongest fractions observed in the LLL take the form $\nu{=}n/(2pn{\pm}1)$, with $n,p$ positive integers, as these corresponds to the $\nu^{*}{=}n$ integer quantum Hall (IQHE) states of CFs. The theory of free CFs also produces states only at odd-denominator fractions.
The experimental observation of a plateau at filling factor $\nu{=}5/2$ in the second Landau level (SLL) [9] showed that FQHE could also arise at even-denominator fractions. Moore and Read [10], using the methods of CFT, postulated a non-Abelian Pfaffian (Pf) wave function to describe this even-denominator state which was subsequently shown to give a good representation of the exact SLL Coulomb ground state at half-filling [11, 12]. Read and Rezayi [13] then generalized the Moore-Read construction to produce a family of states, some of which can capture other experimentally observed plateaux in the SLL [14, 15]. Some experimentally observed fractions in the SLL still lie beyond the purview of the aforementioned theories [16, 17]. In particular, signatures of FQHE have been observed in the zeroth LL (ZLL) of bilayer graphene (BLG), which is believed to stabilize states analogous to those in the SLL, at fractions $6/13$ and $8/17$ and their hole-conjugates $7/13$ and $9/17$ [18, 19, 20, 21] (Throughout this work, states related by particle-hole symmetry would be considered on an equal footing and thus it suffices to just consider $\nu{\leq}1/2$.).
@@
where $\Phi_{n}$ is the Slater determinant state of $n$ filled LLs of electrons with $\Phi_{{-}n}{\equiv}\Phi_{\bar{n}}{=}[\Phi_{n}]^{*}$, and $\mathcal{P}_{\rm LLL}$ implements projection to the LLL as is required in the infinite field limit. Using the fact that the partons are exposed to the same external magnetic field as the electrons and have the density as the electrons, one can show that the charge of the $\alpha$ parton species $q_{\alpha}{=}{-}e\nu/n_{\alpha}$, where $\nu{=}[\sum_{\alpha{=}1}^{l}n^{-1}_{\alpha}]^{-1}$ [7]. The Wen-Zee shift [32] of the parton state of Eq. (1) is $\mathcal{S}{=}\sum_{\alpha{=}1}^{l}n_{\alpha}$ and is thus always an integer. However, an FQHE state can also have a fractional shift [22, 33, 34, 35], and such states (in particular, a family of LH states that include $8/17$ have fractional shifts) are not captured by the wave function of Eq. (1).
@@
The $\bar{1}\bar{2}1^{3}$ is topologically equivalent (and nearly identical) to the $2/3$ Jain state [46] and the latter is known to give a reasonable description of the Coulomb ground state in the entire ZLL of BLG [42]. The $\bar{2}^{2}1^{3}$ is a state that occurs at half-filling and lies in the same universality class as the aPf [28]. Encouragingly, the $\bar{2}^{2}1^{3}$ provides a better representation of the Coulomb ground state in the ZLL of BLG near the SLL Coulomb point than the Pf state [42]. The $n{=}3$ member of the sequence given in Eq. (2) is a candidate state that occurs at $6/13$ where FQHE has been observed both in the SLL [17] and the ZLL of BLG [18, 19, 20, 21]. The $\bar{3}\bar{2}1^{3}$ wave function has been shown to give a good description of the exact SLL Coulomb ground state [48] and in the Supplemental Material (SM) [49], we show that it gives a good description of the ground state in the ZLL of BLG too. It turns out that the $\bar{3}\bar{2}1^{3}$ state lies in the same universality class as the corresponding LH state [48].
@@
where $\Psi_{8/13}{\equiv}\mathcal{P}_{\rm LLL}\Phi^{2}_{1}\Phi_{8/3}$ and the wave function $\Phi_{8/3}$ is constructed by filling the lowest two LLs of electrons and forming the $2/3$ Jain state (or equivalently, the particle-hole conjugate of the $1/3$ Laughlin state) in the third LL. The 8/3 wave function cannot be broken down into a product of IQHE states as it has a fractional shift of 5/2. The 8/17 state of Eq. (3) occurs at shift ${-}3/2$ which is different from the LH state (shift 5/2) and therefore the two states are topologically distinct.
@@
In Fig. 1 we show overlaps of the $\overline{(8/3)}\bar{2}1^{3}$ with the exact Coulomb ground state in the ZLL of BLG for $N{=}12$. The $8/3$ state is constructed by brute-force projection to the LLL and we then use this $8/3$ state to construct the $8/17$ state given in Eq. (3) following the method outlined in Refs. [58, 30] that involves expanding the state in the space of all $L{=}0$ states. The overlaps of the parton state with the exact state are small for all values of $\theta$ in the ZLL of BLG indicating that the parton state does not give a good microscopic description of the experimentally observed $8/17$ state [59, 19, 20, 21]. Nevertheless, the parton state may lie in the same topological phase as the experimentally realized $8/17$ state. For the parton state, the only system accessible to ED is $N{=}12$ since the next system size of $N{=}20$ is currently beyond our reach (Hilbert space dimension ${\simeq}1.43{\times}10^{10}$).
The LH state is not readily amenable to a numerical construction and we have not been able to obtain its wave function. However, we can carry out ED to obtain the ground state for $N{=}20$ (Hilbert space dimension ${\simeq}1.38{\times}10^{9}$) at the shift corresponding to the LH state. In Fig. 2(b) we show the ground state $L$ as a function of $\theta$ for this system. There is a narrow region of parameter space of small to intermediate fields where the ground state at the LH shift in the ZLL of BLG for this system is uniform (Moore-Read quasiholes give a good description of the exact Coulomb ground state here [49].) while the exact SLL Coulomb ground state (that occurs at zero field) at this shift is not uniform. Note that for $N{=}12$ at the LH shift, the ground state has $L{=}0$ irrespective of $\theta$.
@@
If it turns out that the experimentally observed 8/17 state lies in the universality class of $\overline{(8/3)}\bar{2}1^{3}$ state, then it would call into question the idea that looking at the nearby fractions one can tell if the even-denominator state seen at half-filling is in the Pf or the aPf universality class [18, 19, 20, 21, 25]. Finally, we mention the possibility of multi-component states at 8/17 where the different components could represent valley, spin, or orbital degree of freedom. The $\overline{(8/3)}\bar{2}1^{3}$ state readily admits the possibility of unpolarized states building on the partially-polarized and singlet-states at $\nu{=}8/3$ and the singlet states at $\nu{=}2$. It is possible that these states could potentially be relevant for certain interactions. Potentially all possible FQHE states lend themselves to a parton description where the partons themselves form simple IQHE or FQHE states.
@@
Nevertheless, the exact LH state should lie within the Moore-Read quasihole manifold. Thus, we have computed the overlaps of the Moore-Read quasiholes with the exact Coulomb ground state in the zeroth LL of bilayer graphene. Such a calculation can only be performed at $N{=}12$. As mentioned in the main text, the ground state in the ZLL of BLG has $L{=}0$ irrespective of $\theta$. There are exactly 3 Moore-Read quasiholes states with $L{=}0$ at $N{=}12$ and $232l{=}$ 23 flux quanta. While the quasihole overlap calculation is out of reach for $N{=}20$, we have also computed as a proxy the $\langle H_{3}\rangle$ and $\langle\Delta H_{3}\rangle$, where $H_{3}$ is the model 3-body interaction which realizes the Moore-Read Pfaffian state [72], of the exact Coulomb ground state in the zeroth LL of bilayer graphene for $N{=}20$. For the exact LH state, we would get $\langle H_{3}\rangle{=}\langle\Delta H_{3}\rangle{=}0$. These results are shown in Fig. S2. We find that both the $H_{3}$-mean and its standard deviation for the exact Coulomb ground state exhibit a minimum in the region where the ground state is uniform. For completeness, we point out that such calculations are demanding. To obtain the Coulomb ground state of the bilayer graphene for a given $\theta$, the average clock running time is 130h on two nodes with dual AMD EPYC 7542 processors (32 cores per CPU) with 2Tb of RAM each. On the other hand, an expectation value of $H_{3}$ requires on the order 250h using one such node.

### v2

@@ Abstract
The Haldane-Halperin hierarchy postulates that the quasiholes or quasiparticle excitations of a parent FQHE state condense into a daughter FQHE state and this scheme can produce candidate states for any odd-denominator fraction [4, 5]. Jain’s CF theory builds many-particle correlations by the process of vortex attachment wherein electrons bind zeros of the many-body wave function to turn into CFs which to the zeroth-order approximation are taken to be non-interacting [6]. A vast majority of the FQHE phenomenology, especially that observed in the lowest Landau level (LLL), is well-described by the CF theory. In particular, the CF theory explains why the strongest fractions observed in the LLL take the form $\nu{=}n/(2pn{\pm}1)$, with $n,p$ positive integers, as these corresponds to the $\nu^{*}{=}n$ integer quantum Hall effect (IQH) states of CFs. The theory of free CFs also produces states only at odd-denominator fractions.
The experimental observation of a plateau at filling factor $\nu{=}5/2$ in the second Landau level (SLL) [9] showed that FQHE could also arise at even-denominator fractions. Moore and Read [10], using the methods of CFT, postulated a non-Abelian Pfaffian (Pf) wave function to describe this even-denominator state which was subsequently shown to give a good representation of the exact SLL Coulomb ground state at half-filling [11, 12]. Read and Rezayi [13] then generalized the Moore-Read construction to produce a family of states, some of which can capture other experimentally observed plateaux in the SLL [14, 15]. Some experimentally observed fractions in the SLL still lie beyond the purview of the aforementioned theories [16, 17]. In particular, signatures of FQHE have been observed in the zeroth LL (ZLL) of bilayer graphene (BLG), which is believed to stabilize states analogous to those in the SLL, at fractions $6/13$ and $8/17$ and their hole-conjugates $7/13$ and $9/17$ [18, 19, 20, 21, 22] (Throughout this work, states related by particle-hole symmetry would be considered on an equal footing and thus it suffices to just consider $\nu{\leq}1/2$.).
@@
where $\Phi_{n}$ is the Slater determinant state of $n$ filled LLs of electrons with $\Phi_{{-}n}{\equiv}\Phi_{\bar{n}}{=}[\Phi_{n}]^{*}$, and $\mathcal{P}_{\rm LLL}$ implements projection to the LLL as is required in the infinite field limit. Using the fact that the partons are exposed to the same external magnetic field as the electrons and have the same density as the electrons, one can show that the charge of the $\alpha$ parton species $q_{\alpha}{=}{-}e\nu/n_{\alpha}$, where $\nu{=}[\sum_{\alpha{=}1}^{l}n^{-1}_{\alpha}]^{-1}$ [7]. The Wen-Zee shift [33] of the parton state of Eq. (1) is $\mathcal{S}{=}\sum_{\alpha{=}1}^{l}n_{\alpha}$ and is thus always an integer. However, an FQHE state can also have a fractional shift [23, 34, 35, 36], and such states (in particular, a family of LH states that include $8/17$ have fractional shifts) are not captured by the wave function of Eq. (1).
@@
The $\bar{1}\bar{2}1^{3}$ is topologically equivalent (and nearly identical) to the $2/3$ Jain state [47] and the latter is known to give a reasonable description of the Coulomb ground state in the entire ZLL of BLG [43]. The $\bar{2}^{2}1^{3}$ state occurs at half-filling and lies in the same universality class as the aPf [29]. Encouragingly, the $\bar{2}^{2}1^{3}$ provides a better representation of the Coulomb ground state in the ZLL of BLG near the SLL Coulomb point than the Pf state [43]. The $n{=}3$ member of the sequence given in Eq. (2) is a candidate state that occurs at $6/13$ where FQHE has been observed both in the SLL [17] and the ZLL of BLG [18, 19, 20, 21, 22]. The $\bar{3}\bar{2}1^{3}$ wave function has been shown to give a good description of the exact SLL Coulomb ground state [49] and in the Supplemental Material (SM) [50], we show that it gives a good description of the ground state in the ZLL of BLG too. It turns out that the $\bar{3}\bar{2}1^{3}$ state lies in the same universality class as the corresponding LH state [49].
@@
where $\Psi_{8/13}{\equiv}\mathcal{P}_{\rm LLL}\Phi^{2}_{1}[\Phi_{8/3}]^{*}$ and the wave function $\Phi_{8/3}$ is constructed by filling the lowest two LLs of electrons and forming the $2/3$ Jain state (or equivalently, the particle-hole conjugate of the $1/3$ Laughlin state) in the third LL. The 8/3 wave function cannot be broken down into a product of IQHE states as it has a fractional shift of 5/2. The 8/17 state of Eq. (3) occurs at shift ${-}3/2$ which is different from the LH state (shift 5/2) and therefore the two states are topologically distinct.
@@
In Fig. 1 we show overlaps of the $\overline{(8/3)}\bar{2}1^{3}$ with the exact Coulomb ground state in the ZLL of BLG for $N{=}12$. The $8/3$ state is constructed by brute-force projection to the LLL and we then use this $8/3$ state to construct the $8/17$ state given in Eq. (3) following the method outlined in Refs. [59, 31] that involves expanding the state in the space of all $L{=}0$ states. The overlaps of the parton state with the exact state are small for all values of $\theta$ in the ZLL of BLG indicating that the parton state does not give a good microscopic description of the experimentally observed $8/17$ state [60, 19, 20, 21, 22]. Nevertheless, the parton state may lie in the same topological phase as the experimentally realized $8/17$ state. For the parton state, the only system accessible to ED is $N{=}12$ since the next system size of $N{=}20$ is currently beyond our reach (Hilbert space dimension ${\simeq}1.43{\times}10^{10}$).
The LH state is not readily amenable to a numerical construction and we have not been able to obtain its wave function. However, we can carry out ED to obtain the ground state for $N{=}20$ (Hilbert space dimension ${\simeq}1.38{\times}10^{9}$) at the shift corresponding to the LH state. In Fig. 2(b) we show the ground state $L$ as a function of $\theta$ for this system. There is a narrow region of parameter space of small to intermediate fields where the ground state at the LH shift in the ZLL of BLG for this system is uniform (Moore-Read quasiholes give a good description of the exact Coulomb ground state here [50].) while the exact SLL Coulomb ground state (that occurs at zero field) at this shift is not uniform. Note that for $N{=}12$ at the LH shift, the ground state has $L{=}0$ for all $\theta$.
@@
If it turns out that the experimentally observed 8/17 state lies in the universality class of $\overline{(8/3)}\bar{2}1^{3}$ state, then it would call into question the idea that by looking at the nearby fractions one can tell if the even-denominator state seen at half-filling is in the Pf or the aPf universality class [18, 19, 20, 21, 26, 71, 72, 22]. Finally, we mention the possibility of multi-component states at 8/17 where the different components could represent valley, spin, or orbital degree of freedom. The $\overline{(8/3)}\bar{2}1^{3}$ state readily admits the possibility of unpolarized states building on the partially-polarized and singlet-states at $\nu{=}8/3$ and the singlet state at $\nu{=}2$. It is possible that these states could potentially be relevant for certain interactions. Potentially all possible FQHE states lend themselves to a parton description where the partons themselves form simple IQHE or FQHE states.
@@
Nevertheless, the exact LH state should lie within the Moore-Read quasihole manifold. Thus, we have computed the overlaps of the Moore-Read quasiholes with the exact Coulomb ground state in the zeroth LL of bilayer graphene. Such a calculation can only be performed at $N{=}12$. As mentioned in the main text, the ground state in the ZLL of BLG has $L{=}0$ for all $\theta$. There are exactly 3 Moore-Read quasiholes states with $L{=}0$ at $N{=}12$ and $232l{=}$ 23 flux quanta. While the quasihole overlap calculation is out of reach for $N{=}20$, we have also computed as a proxy the $\langle H_{3}\rangle$ and $\langle\Delta H_{3}\rangle$, where $H_{3}$ is the model 3-body interaction which realizes the Moore-Read Pfaffian state [74], of the exact Coulomb ground state in the zeroth LL of bilayer graphene for $N{=}20$. For the exact LH state, we would get $\langle H_{3}\rangle{=}\langle\Delta H_{3}\rangle{=}0$. These results are shown in Fig. S2. We find that both the $H_{3}$-mean and its standard deviation for the exact Coulomb ground state exhibit a minimum in the region where the ground state is uniform. For completeness, we point out that such calculations are demanding. To obtain the Coulomb ground state of the bilayer graphene for a given $\theta$, the average clock running time is 130h on two nodes with dual AMD EPYC 7542 processors (32 cores per CPU) with 2Tb of RAM each. On the other hand, an expectation value of $H_{3}$ requires on the order 250h using one such node.